OpenStudy (anonymous):

WILL GIVE MEDAL AND BE YOUR FAN!!!! What value of Z makes the equation true? 5/6= z/12 A. 15 B. 5 C. 10 D. 6 I think it's C

OpenStudy (anonymous):

\(\Large \frac{5}{6}=\frac{z}{12} \Rightarrow z=\frac{12\times5}{6}=10\)

OpenStudy (anonymous):

thank you soo much!!

OpenStudy (anonymous):

^^ Welcome